OpenStudy (anonymous):

help probability question! Dillon has a bag of marbles. There are 3 blue, 2 green, 2 white, 3 black, and some red marbles in the bag. If the probability of pulling out a red marble is 3/4, how many red marbles are in the bag? i think its 20 but im not sure A. 30 B. 40 C. 20 D. 50

OpenStudy (cwrw238):

there are 10 marbles in the bag which are not red for P(red) to be 3/4 there must be 30 red marbles (total marbles = 40) P(red) = 30/40 = 3/4

OpenStudy (anonymous):

3+2+2+3 = 10 known marbles in the bag. If the probability of picking a red is 3/4, then 75% of the marbles are red, 25% of them are other colors. If 25% marbles aren't red, multiply the 10 by 4, and there are 40 total marbles, 30 of which are red. B.